how many milliliters of a 25% acid solution and how many milliliters of a 6 % acid solution must be mixed yo get 285 milliliters of a 12%acid solution
Answer by addingup(3677)

Let the amount at 25% be x
Then the amount at 6% will be 285-x
:
0.25x+0.06(285-x) = 0.12(285)
0.25x+17.1-0.06x = 34.2
0.19x = 17.1
x = 90ml of the 25% solution
mixed with
285-90 = 195ml of the 6% solution